OpenStudy (anonymous):

help me please!!!! y= -2/3x +400 Describe how you would graph this line using the slope-intercept method.

OpenStudy (anonymous):

It's best to start off with the two axis intercepts. You already know one just by reading this equation, the y-axis.

OpenStudy (anonymous):

The y axis is (0, 400)

OpenStudy (anonymous):

so the slope is rise over run and 400 is the y intercept.

OpenStudy (anonymous):

Yes, and solve for the x-axis by settings y = 0 = -2/3x + 400

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 } x = 400\]

OpenStudy (anonymous):

So the x-axis is (600, 0)

OpenStudy (anonymous):

Connect your x-axis with your y-axis, and extend both ends of the line to the edge of your graph. You're done!

OpenStudy (anonymous):

3 Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences.

OpenStudy (anonymous):

@math1234

OpenStudy (anonymous):

Functional notation means that you express y in terms a function of x.

OpenStudy (anonymous):

\[f(x) = -\frac{ 2 }{ 3 } x + 400\]

OpenStudy (anonymous):

 start at 400 go down two over three